# TrialBench: Multi-Modal Artificial Intelligence-Ready Clinical Trial Datasets

### v1

@@ 4.1 Multi-modal Deep Neural Networks
For all the classification and regression tasks, we apply state-of-the-art multi-modal deep neural networks to represent multi-modal features. Each representation is an embedding vector with continuous values. Then, we combine these representations and make the prediction. For the eligibility criteria design task, we use OpenAI ChatGPT APIiiihttps://openai.com/index/openai-api/ with the prompt to produce eligibility criteria.
Categorical and Numerical Features
@@
[…]
@@
After obtaining the representations of multi-modal data, we concatenate these representations, feed the concatenated vector into MLP, and make the prediction. For binary classification tasks (e.g., trial approval prediction), we use sigmoid function as the activation function in the output layer to yield predicted probability; for multi-category classification tasks (e.g., trial failure reason identification), we use softmax as the activation function in the output layer to produce probability distribution over all the categories; for regression tasks (e.g., trial duration prediction), we do not use activation function in the output layer to produce continuous-valued prediction. We use cross-entropy criterion as the loss function for classification tasks and mean-square error (MSE) as the loss function for regression tasks.
@@ 4.2 Experimental Setup
All the code is implemented in Python 3.8. All the deep learning models are implemented in PyTorch. We use GPT 4.0 for data annotation and some generation tasks. The total cost was around \$100 US dollars. The embedding size of all the representations is set to 100. We use Adam [60] as the numerical optimizer to minimize the loss function with an initial learning rate at 1e31e{-3} and zero weight decay. The batch size is set to 64. The maximal training epochs is set to 20.
@@
For classification tasks, we assess the model performance using accuracy, PR-AUC (the area under the Precision-Recall curve), F1 score (the harmonic mean of precision and recall), and ROC-AUC (the Area Under the Receiver Operating Characteristic Curve). For regression tasks, we use RMSE (Root Mean Squared Error), MAE (Mean Absolute Error), Concordance Index, and Pearson Correlation as metrics. For generation tasks (eligibility criteria design), we design some semantic metrics to measure the alignment between real and designed criteria, including text embeddings’ cosine similarity, informativeness, and redundancy, detailed in supplementary materials.
4.3 Experimental Results
@@ 5 Usage Note
Code Availability
The curated dataset and relevant code are publicly available at https://github.com/ML2Health/ML2ClinicalTrials/tree/main/AI4Trial.
@@
Figure 7: Input Variables for trial duration prediction.
Figure 9: Feature for trial approval prediction.
Figure 10: Feature for trial failure reason identification.
Figure 11: Feature for eligibility criteria design.
@@ Appendix A Evaluation Metrics
  • F1. The F1 score is the harmonic mean of the precision and recall. That is, $\text{F1}=\frac{2}{\frac{1}{\text{precision }}+\frac{1}{\text{recall}}}$.
@@
In the generation tasks (eligibility criteria design, trial summarization), we design a couple of metrics to evaluate the quality of the generated text content.
@@
  • Informativeness. In text generation, informativeness is one of the key metrics for evaluating the quality of generated text. Informativeness measures the richness of useful information provided by the generated text. High informativeness usually means the text contains more relevant details, facts, and insights, rather than simply repeating known information or containing empty content. We use TF-IDF (Term Frequency-Inverse Document Frequency) to measure the informativeness of the generated text relative to a reference text: (1) TF-IDF vectorization: Combine the generated text and reference text, and vectorize them using the TF-IDF method. TF-IDF is a statistical method used to evaluate the importance of a word in a document. (2) cosine similarity: Calculate the cosine similarity between the generated text’s TF-IDF vector and the reference text’s average TF-IDF vector. The higher the cosine similarity, the more similar the generated text is to the reference text. (3) informativeness score: The informativeness score is calculated as 1 - cosine similarity. A higher score indicates that the generated text contains more different information relative to the reference text, thus having greater informativeness.

### v2

@@ 4.1 Multi-modal Deep Neural Networks
For all the classification and regression tasks, we apply various deep neural networks to represent multi-modal features. Each representation is an embedding vector with continuous values. Then, we concatenate these representations, feed them into multiple layer perceptron (MLP), and make the prediction. For the eligibility criteria design task, we use OpenAI ChatGPT APIiiihttps://openai.com/index/openai-api/ with the prompt to produce eligibility criteria.
Categorical and Numerical Tabular Features
@@
$\mathbf{m}_{uv}^{(l)}=f_{1}\bigg{(}\mathbf{e}_{u}\oplus\mathbf{e}_{uv}^{(l-1)}\oplus\sum_{w\in\mathcal{N}(u)\backslash v}\mathbf{m}_{wu}^{(l-1)}\bigg{)},\ \ \ \ l=1,\cdots,L,$ (3)
@@
After obtaining the representations of multi-modal data, we concatenate these representations, feed the concatenated vector into multiple-layer perceptron (MLP), and make the prediction. For binary classification tasks (e.g., trial approval prediction), we use sigmoid function as the activation function in the output layer to yield predicted probability; for multi-category classification tasks (e.g., trial failure reason identification), we use softmax as the activation function in the output layer to produce probability distribution over all the categories; for regression tasks (e.g., trial duration prediction), we do not use activation function in the output layer to produce continuous-valued prediction. We use cross-entropy criterion as the loss function for classification tasks and mean-square error (MSE) as the loss function for regression tasks.
@@ 4.2 Experimental Setup
All the code is implemented in Python 3.8. All the deep learning models are implemented in PyTorch, and we use GPT 4.0 for data annotation and the generation tasks. The embedding size of all the representations is set to 100. We use Adam [60] as the numerical optimizer to minimize the loss function with an initial learning rate at $1e{-3}$ and zero weight decay. The batch size is set to 64. The maximal training epochs is set to 20.
@@
For classification tasks, we assess the model performance using accuracy, PR-AUC (the area under the Precision-Recall curve), F1 score (the harmonic mean of precision and recall), and ROC-AUC (the Area Under the Receiver Operating Characteristic Curve). For regression tasks, we use RMSE (Root Mean Squared Error), MAE (Mean Absolute Error), Concordance Index, and Pearson Correlation as metrics. For generation tasks (eligibility criteria design), we design some semantic metrics to measure the alignment between real and designed criteria, including text embeddings’ cosine similarity, informativeness, and redundancy, detailed in the Appendix.
4.3 Validation Results
@@ 5 Usage Note
Code & Data Availability
The curated dataset are available at https://huggingface.co/datasets/ML2Healthcare/ClinicalTrialDataset and the validation frameworks are available at https://github.com/ML2Health/ML2ClinicalTrials/tree/main/AI4Trial.
@@ Appendix A Evaluation Metrics
[…]
@@
In the generation tasks (e.g., eligibility criteria design), we design a couple of metrics to evaluate the quality of the generated text content.
@@
  • Informativeness. In text generation, informativeness is one of the key metrics for evaluating the quality of generated text. Informativeness measures the richness of useful information provided by the generated text. High informativeness usually means the text contains more relevant details, facts, and insights, rather than simply repeating known information or containing empty content. We use TF-IDF (Term Frequency-Inverse Document Frequency, a well-known metric in natural language processing) to measure the informativeness of the generated text relative to a reference text: (1) TF-IDF vectorization: Combine the generated text and reference text, and vectorize them using the TF-IDF method. TF-IDF is a statistical method used to evaluate the importance of a word in a document. (2) cosine similarity: Calculate the cosine similarity between the generated text’s TF-IDF vector and the reference text’s average TF-IDF vector. The higher the cosine similarity, the more similar the generated text is to the reference text. (3) informativeness score: The informativeness score is calculated as 1 minus cosine similarity. A higher score indicates that the generated text contains more different information relative to the reference text, thus having greater informativeness.
@@
Figure 6: Input Variables for trial duration prediction.
Figure 8: Feature for trial approval prediction.
Figure 9: Feature for trial failure reason identification.
Figure 10: Feature for eligibility criteria design.